# Extreme mass distributions for $k$-increasing quasi-copulas

### v1

@@ Abstract.
The rating of quasi-copula problems in the dependence modeling community has recently raised in spite of the lack of probability interpretation of quasi-copulas. The trendsetting paper [3] proposes the study of $k$-increasing property for some $k\leq d$ as a property of $d$-variate quasi-copulas that would shed some light on what is in-between. This hierarchy of classes extends the bivariate notion of supermodularity property. The same authors propose a number of open problems in the continuation of this paper [4]. Their Open problem 5 asks for the extreme values of the mass distributions associated with multivariate quasi-copulas and was recently solved by the authors of this paper [13]. The main goal of the present paper is to solve the maximal-volume problem (in absolute value) within each of the previously mentioned subclasses. By formulating and solving suitably simplified primal and dual linear programs, we derive the exact maximal negative and positive masses together with the corresponding extremal boxes.
@@ 1. Introduction
The seminal paper [4] by Arias-García, Mesiar, and De Baets (2020) presents a survey of the most significant results and open problems in quasi-copula theory. The fifth problem on their list—also referred to as Hitchhiker’s problem #5—asks for the maximal negative and maximal positive mass over all boxes and over all quasi-copulas, and for the characterization of the boxes attaining these extrema. Here, maximal negative refers to the negative value with the largest absolute magnitude.
This problem has now been solved in full generality using a linear programming approach in our recent work [13]. Special cases had previously been addressed in [12, 6, 14, 5]. It turns out that the maximal volumes grow exponentially, and the extremal values can be computed using a simple algorithm determining the smallest positive integer satisfying a certain inequality.
@@
In [8, 9], the authors studied properties of 2-increasing binary aggregation operators and provided a method for constructing copulas from them. In dimensions $d>2$, beyond the class of 2-increasing functions, there exists a hierarchy of intermediate classes satisfying the $k$-increasing property for $k=2,\ldots,d$. These classes were introduced in [3, Section 5] (see also [4, Section 6.3]), extending the volume-based characterization of quasi-copulas to $k$-dimensional slices of the $d$-dimensional box. In this way one obtains special subclasses of quasi-copulas: they are strictly contained within the class of all quasi-copulas (corresponding to the 1-increasing property) and contain all copulas (corresponding to the $d$-increasing property).
The motivation for the present paper is to study the maximal volume problem described above within these intermediate classes of quasi-copulas, and thereby to quantify how restrictive the $k$-increasing property becomes in terms of maximal volumes. Our results may also serve as indicators of the “distance” between the classes of quasi-copulas (resp. copulas) and the class of $k$-increasing quasi-copulas. As noted in [3, Section 6], certain properties of bivariate copulas do not extend to higher dimensions $d$, but they do hold for some of the intermediate classes associated with small $k$, suggesting that quasi-copulas are more closely related to these. On the other hand, classes with larger $k$ have been used to study copulas, illustrating the close connection between the two settings.
In the bivariate case, the 2-increasing property coincides with supermodularity (see, e.g., [11, Definition 2.1]). Consequently, the notion of $k$-increasingness can be viewed as a natural generalization of supermodularity to higher dimensions. Very recently, a related problem to the main problem of this paper, was studied by Anzilli and Durante [2], who derived bounds on the average $F$-volume of closed rectangles in $0,1]^{2}$ for a supermodular aggregation function $F$. These average-volume bounds offer an additional perspective for understanding such aggregation functions and, in particular, copulas.
@@
The paper is organized as follows. In Section 2 we present our main results (Theorems 1 and 2), together with a numerical analysis of the extremal values (Tables 1 and 2). Section 3 formulates the maximal-volume problems as linear programs (Proposition 1), simplifies them using a symmetrization argument (Corollary 1), introduces new sets of variables (Proposition 3), and derives the dual linear programs (Proposition 4). In Section 5 we prove Theorems 1 and 2 by solving these dual programs exactly. Finally, in Section 6 we summarize the paper and give some directions for further work.
@@ 2. Statements of the main theorems
In this section we introduce the notation and state our main results, i.e., Theorem 1 solves the maximal negative volume problem for $k$-increasing quasi-copulas, while Theorem 2 solves the maximal positive volume problem. Concrete numerical solutions examples are also presented in the form of tables. (See Tables 1 and 2 for the minimal and the maximal volume question, respectively.)
@@
If $\mathcal{D}=[0,1]^{d}$ and $F$ satisfies the boundary, the monotonicity and the Lipschitz conditions, then $F$ is called a $d$-variate quasi-copula (or $d$-quasi-copula). We will omit the dimension $d$ when it is clear from the context and write quasi-copula for short.
Let ${\mathcal{B}}:=\prod_{j=1}^{d}[x_{j},y_{j}]$, where $0\leq x_{j}\leq y_{j}\leq 1$ for each $j$. Next we recall the definition of a $k$-dimensional section of a function $F:{\mathcal{B}}\to[0,1]$ (see [3, Definition 5]). For any $\underline{a}:=(a_{1},\ldots,a_{d})\in{\mathcal{B}}$ and a set of indices $A\subseteq\{1,2,\ldots,d\}$ with $0<\operatorname{card}A=k\leq d$, the $k$-dimensional slice (or $k$-slice) ${\mathcal{B}}_{\underline{a},A}$ of $\mathcal{B}$} with fixed values given by $\underline{a}$ in the positions determined by $A$, is defined by
@@
The solution to the maximal negative volume problem, i.e., the largest in absolute value among negative ones, over the set of $k$-increasing $d$-quasi-copulas, is the following.
@@
Assume the notation above. Let $d,k\in\mathbb{N}$ and $2\leq k\leq d$. The maximal negative volume $V_{Q}({\mathcal{B}})$ of some $d$-box $\mathcal{B}$} over all $k$-increasing $d$-quasi-copulas $Q$ is equal to
$w_{k,d,-}:=\min\Big(0,\min_{i=k,\ldots,d}(-1)^{d-i}\binom{d-k}{i-k}\left(\gamma_{i}^{(k)}\right)^{-1}\Big),$
@@
$w_{k,d,-}=(-1)^{d-i_{0}}\binom{d-k}{i_{0}-k}\left(\gamma_{i_{0}}^{(k)}\right)^{-1}.$
One of the realizations of $Q$ and $\mathcal{B}$} such that $V_{Q}({\mathcal{B}})=w_{k,d,-}$ has the following properties:
@@
The solution to the maximal positive volume question over the set of $k$-increasing $d$-quasi-copulas, is the following.
@@
Assume the notation above. Let $d,k\in\mathbb{N}$ and $2\leq k\leq d$. The maximal positive volume $V_{Q}({\mathcal{B}})$ of some $d$-box $\mathcal{B}$} over all $k$-increasing $d$-quasi-copulas $Q$ is equal to
$w_{k,d,+}:=\max_{i=k,\ldots,d}(-1)^{d-i}\binom{d-k}{i-k}(\gamma_{i}^{(k)})^{-1},$
@@
In Table 1 and 2111The numerical analysis in arithmetic over $\mathbb{Q}$ was performed using the software tool Mathematica [15]. The source code is available at https://github.com/ZalarA/Quasi-copulas-k-increasing-extreme-volumes. we give maximal negative and positive volumes of the $d$-boxes over all $k$-increasing $d$-quasi-copulas for dimensions up to 15. The cases $k\geq 2$ are computed according to Theorems 1 and 2 above, while the case $k=1$ was studied in our previous work [13, Theorems 1 and 2].
Table 1. Minimal values of $V_{Q}({\mathcal{B}})$ over all $d$-variate $k$-increasing quasi-copulas $Q$ and all $d$-boxes ${\mathcal{B}}\subseteq[0,1]^{d}$.
@@
Table 2. Maximal values of $V_{Q}({\mathcal{B}})$ over all $d$-variate $k$-increasing quasi-copulas $Q$ and all $d$-boxes ${\mathcal{B}}\subseteq[0,1]^{d}$.
@@ 3. Towards the proof of Theorems 1 and 2
then the optimal value of (3.2) is the maximal negative volume of some box over all $k$-increasing $d$-quasi-copulas.
Moreover, if there exists an optimal solution to (3.2) where $\min$ is replaced with $\max$, which satisfies (3.3), then the optimal value of (3.2) is the maximal positive volume of some box over all $k$-increasing $d$-quasi-copulas.
@@
The linear program (3.2) is obtained by extending the linear program for general quasi-copulas introduced in [13, Proposition 2] with additional constraints enforcing the $k$-increasing property. By [3, Lemma 2], any $k$-increasing quasi-copula is also $j$-increasing for every $j\leq k$. Nevertheless, in the linear program it is necessary to impose the $j$-increasing constraints explicitly for all $j\leq k$. Otherwise, the resulting quasi-copula would satisfy the $k$-increasing property only on the box $\prod_{i=1}^{k}[a_{i},b_{i}]$, which is insufficient to invoke [3, Lemma 2].
@@
By the same proof as in [5, Theorem 2.1], it follows that $Q$ is a quasi-copula. It remains to prove that it is $k$-increasing. By Lemma 3.4, it suffices to prove that $Q$ is $k$-increasing on each $k$-face of each box ${\mathcal{B}}_{\mathbb{I}}$ for $\mathbb{I}\in\{0,1\}^{d}$. By construction, this is true for $\mathbb{I}_{\text{main}}=(1,1,\ldots,1)$. Let us now consider $\mathbb{I}\in\{0,1\}^{d}\setminus\mathbb{I}_{\text{main}}$. If the face has constant 0 on some component, then $Q$ is constantly equal to 0 on the whole face and $Q$ is $k$-increasing. Otherwise every constant component is equal to $a_{j}$ or $1$. Further on, we can assume that all non-constant components are $a_{j},1]$ and not $0,a_{j}]$, because otherwise we can further project the section on the $j$-th component to $\{a_{j}\}$, hence obtaining a lower dimensional face of ${\mathcal{B}}_{\mathbb{I}_{\text{main}}}$. However, this reduction give us a $k$-dimensional face of ${\mathcal{B}}_{\mathbb{I}_{\text{main}}}$, whence it is $k$-increasing by construction. ∎
@@
The proof is the same as for [5, Proposition 3], but we include it for the sake of completeness. Let $S_{d}$ be the set of all permutations of a $d$-element set $\{1,\ldots,d\}$. For $\Phi\in S_{d}$ and $\mathbb{I}^{(j)}:=(\mathbb{I}_{1}^{(j)},\ldots,\mathbb{I}_{d}^{(j)})\in\{0,1\}^{d}$, let $\Phi(\mathbb{I}^{(j)}):=\big(\mathbb{I}_{\Phi(1)}^{(j)},\ldots,\mathbb{I}_{\Phi(d)}^{(j)}\big)$. For every optimal solution $\big(a_{1}^{\ast},\ldots,a_{d}^{\ast},b_{1}^{\ast},\ldots,b_{d}^{\ast},q_{\mathbb{I}^{(1)}}^{\ast},\ldots,q_{\mathbb{I}^{(2^{d})}}^{\ast}\big)$ to (3.2), $\big(a_{1}^{\ast},\ldots,a_{d}^{\ast},b_{1}^{\ast},\ldots,b_{d}^{\ast},q_{\mathbb{I}^{(1)}}^{\ast},\ldots,q_{\mathbb{I}^{(2^{d})}}^{\ast}\big)$ is also an optimal solution to (3.2), whence an optimal solution to (3.2) of the form (3.7) is equal to $\frac{1}{d!}\sum_{\Phi\in S_{d}}\big(a_{\Phi(1)}^{\ast},\ldots,a_{\Phi(d)}^{\ast},b_{\Phi(1)}^{\ast},\ldots,b_{\Phi(d)}^{\ast},q_{\Phi(\mathbb{I}^{(1)})}^{\ast},\ldots,q_{\Phi(\mathbb{I}^{(2^{d})})}^{\ast}\big).$
@@
We start by $k=2$. By [13, Proposition 5], (3.8) is equivalent to

### v2

# Maximal signed volume for (multivariate) supermodular quasi-copulas

@@ Abstract.
Copulas are the main tool of dependence modeling in statistics and quasi-copulas are their necessary companions. The latter appear, say, as infima or suprema of sets of copulas, they form a huge class and have some unpleasant properties. Their statistical interpretation is challenged by the fact that they may lead to negative volumes of some boxes. So, numerous applications call for an intermediate class and supermodular quasi-copulas is one of them having many useful properties. An excellent measure, Average Rectangular Volume (ARV in short), to clarify and position this class was proposed in the seminal paper by Anzilli and Durante, The average rectangular volume induced by supermodular aggregation functions, J. Math. Anal. Appl. 555 (2026) 21 pp. While supermodularity is a bivariate notion, its extension to $d$-variate case for $d>2$ was recently provided in a key paper by Arias-García, Mesiar, and De Baets, The unwalked path between quasi-copulas and copulas: Stepping stones in higher dimensions, Int. J. of Appr. Reasoning, 80 (2017) pp. 89–99. Here an alternative method to ARV is presented extendable to the multivariate case based on Maximal (in absolute value) Negative Volumes (MNV in short) on boxes thus helping practitioners when seeking the right (quasi-)copula for their problem. Observe that these volumes on copulas are zero, while their values on quasi-copulas, depending on $d$, has been a long-open problem solved only recently. We present a nontrivial extension of this solution to exhibit the main goal of this paper, a measure that clarifies and positions the classes considered based on MNV.
@@ 1. Introduction
Copula [1] is a joint distribution with uniform margins, but when Sklar [2] discovered in 1959 that we could get any distribution on earth when we inserted appropriate univariate distributions as margins into a copula, it became a universal device to solve dependence problems. There are many applications, though, where quasi-copulas are also needed, they form a pretty big class and have some unpleasant properties. They appear, say, as infima or suprema of sets of copulas, but their statistical interpretation is challenged by the fact that they may lead to negative volumes of some boxes. So, numerous applications call for an intermediate class and supermodular quasi-copulas is one of them having many useful properties.
An excellent method, Average Rectangular Volume (ARV in short), to aggregate the knowledge contained in a quasi-copula of this class was proposed by Anzilli and Durante in their seminal paper [3] in 2026. While supermodularity is a bivariate notion [4], its extension to $d$-variate case for $d>2$ was recently provided in a key paper by Arias-García et al. [5] in 2017 and we want to find a parallel notion to ARV for these classes between the classes of quasi-copulas they are contained in and copulas they contain. There seems to be no obvious extension of the method presented in [3], so, we develop a novel method based on maximal (in absolute value) negative volume on boxes. Observe that this value on copulas is zero, while its value on quasi-copulas, depending on $d$, has been a long-open problem solved only recently. Namely, the seminal paper [6] by Arias-García et al. which is in some sense a 2020 sequel of [5] presents a survey of the most significant results and open problems in quasi-copula theory. Their Open Problem #5 asks for the maximal negative and maximal positive mass over all boxes and over all quasi-copulas, for the characterization of the boxes attaining these extremes, and for characterization of quasi-copulas with this property. Here, maximal negative refers to the negative volume with the largest absolute magnitude. The problem seems to have been first proposed in 2002 by Nelsen et al. [7], the trivariate solution was first presented in 2007 by De Beats et al. [8] and the problem was solved in papers [9] in 2023 and [10, 11] in 2025/26.
@@
The notion of supermodularity (see, e.g., [12, Definition 2.1]) coincides with the 2-increasing property in the bivariate case. In dimensions $d\geq\emph{}2$, beyond the class of 2-increasing functions, there exists a hierarchy of intermediate classes satisfying the $k$-increasing property for $k=2,\ldots,d$. These classes were introduced in [5, Section 5] (see also [6, Section 6.3]), extending the volume-based characterization of quasi-copulas to $k$-dimensional slices of the $d$-dimensional box. In this way one obtains special subclasses of quasi-copulas: they are strictly contained within the class of all quasi-copulas (corresponding to the 1-increasing property) and contain all copulas (corresponding to the $d$-increasing property). Consequently, the notion of $k$-increasingness can be viewed as a natural generalization of supermodularity to higher dimensions as we mentioned earlier. The Maximal (in absolute value) Negative Volumes (MNV in short) on boxes and their counterpart Maximal Positive Volumes (MPV in short) bring us to their tie Maximal Signed Volumes (MSV for short), thus providing a possible parallel of ARV.
@@
The manuscript is organized as follows. In Section 2 we present our main results (Theorems 1 and 2), together with a numerical analysis of the extremal values (Tables 1 and 2). Section 3 formulates the maximal-volume problems as linear programs (Proposition 1), simplifies them using a symmetrization argument (Corollary 1), introduces new sets of variables (Proposition 3), and derives the dual linear programs (Proposition 4). In Section 4 we prove Theorems 1 and 2 by solving these dual programs exactly. Finally, in Section 5 we summarize the paper and give some directions for further work.
@@ 2. Statements of the main theorems
In this section we introduce the notation and state our main results, i.e., Theorem 1 solves the maximal negative volume problem for $k$-increasing quasi-copulas, while Theorem 2 solves the maximal positive volume problem. Concrete numerical solution examples are also presented in the form of tables. (See Tables 1 and 2 for the minimal and the maximal volume question, respectively.)
@@
If $\mathcal{D}=[0,1]^{d}$ and $F$ satisfies the boundary, the monotonicity and the Lipschitz conditions, then $F$ is called a $d$-variate quasi-copula. We will omit the dimension $d$ when it is clear from the context and write quasi-copula for short.
Let ${\mathcal{B}}:=\prod_{j=1}^{d}[x_{j},y_{j}]$, where $0\leq x_{j}\leq y_{j}\leq 1$ for each $j$. Next we recall the definition of a $k$-dimensional section of a function $F:{\mathcal{B}}\to[0,1]$ (see [5, Definition 5]). For any $\underline{a}:=(a_{1},\ldots,a_{d})\in{\mathcal{B}}$ and a set of indices $A\subseteq\{1,2,\ldots,d\}$ with $0<\operatorname{card}A=k\leq d$, the $k$-dimensional slice (or $k$-slice) ${\mathcal{B}}_{\underline{a},A}$ of $\mathcal{B}$} with fixed values given by $\underline{a}$ in the positions determined by $A$, is defined by
@@
We denote by $\mathrm{MNV}(d,k)$ the maximal negative volume $V_{Q}({\mathcal{B}})$, i.e., the largest in absolute value among negative ones, of some $d$-box $\mathcal{B}$} over all $k$-increasing $d$-variate quasi-copulas $Q$. The following result provides an explicit formula for $\mathrm{MNV}(d,k)$.
@@
Assume the notation above. Let $d,k\in\mathbb{N}$ and $2\leq k\leq d$. Then
$\mathrm{MNV}(d,k)=\min\Big(0,\min_{i=k,\ldots,d}(-1)^{d-i}\binom{d-k}{i-k}\left(\gamma_{i}^{(k)}\right)^{-1}\Big),$
@@
$\mathrm{MNV}(d,k)=(-1)^{d-i_{0}}\binom{d-k}{i_{0}-k}\left(\gamma_{i_{0}}^{(k)}\right)^{-1}.$
One of the realizations of $Q$ and $\mathcal{B}$} such that $V_{Q}({\mathcal{B}})=\mathrm{MNV}(d,k)$ has the following properties:
@@
We denote by $\mathrm{MPV}(d,k)$ the maximal positive volume $V_{Q}({\mathcal{B}})$ of some $d$-box $\mathcal{B}$} over all $k$-increasing $d$-variate quasi-copulas $Q$. The following result provides an explicit formula for $\mathrm{MPV}(d,k)$.
@@
Assume the notation above. Let $d,k\in\mathbb{N}$ and $2\leq k\leq d$. Then
$\mathrm{MPV}(d,k)=\max_{i=k,\ldots,d}(-1)^{d-i}\binom{d-k}{i-k}(\gamma_{i}^{(k)})^{-1},$
@@
We may also consider the maximal signed volume as the pair $\mathrm{MSV}(d,k)=(-\mathrm{MNV}(d,k),\mathrm{MPV}(d,k))$; observe that $\mathrm{MSV}(2,2)$ may be seen as a possible statistical counterpart to ARV of [3].
In Table 1 and 2111The numerical analysis in arithmetic over $\mathbb{Q}$ was performed using the software tool Mathematica [13]. The source code is available at https://github.com/ZalarA/Quasi-copulas-k-increasing-extreme-volumes. we give the values of $\mathrm{MNV}(d,k)$ and $\mathrm{MPV}(d,k)$ for dimensions up $d$ to 15. The cases $k\geq 2$ are computed according to Theorems 1 and 2 above, while the case $k=1$ was studied in our previous work [11, Theorems 1 and 2].
Table 1. Explicit values of $\mathrm{MNV}(d,k)$ for $2\leq d\leq 15$ and $1\leq k\leq d$.
@@
Table 2. Explicit values of $\mathrm{MPV}(d,k)$ for $2\leq d\leq 15$ and $1\leq k\leq d$.
@@ 3. Towards the proof of Theorems 1 and 2
then the optimal value of (3.2) is the maximal negative volume of some box over all $k$-increasing $d$-variate quasi-copulas.
Moreover, if there exists an optimal solution to (3.2) where $\min$ is replaced with $\max$, which satisfies (3.3), then the optimal value of (3.2) is the maximal positive volume of some box over all $k$-increasing $d$-variate quasi-copulas.
@@
The linear program (3.2) is obtained by extending the linear program for general quasi-copulas introduced in [11, Proposition 2] with additional constraints enforcing the $k$-increasing property. By [5, Lemma 2], any $k$-increasing quasi-copula is also $j$-increasing for every $j\leq k$. Nevertheless, in the linear program it is necessary to impose the $j$-increasing constraints explicitly for all $j\leq k$. Otherwise, the resulting quasi-copula would satisfy the $k$-increasing property only on the box $\prod_{i=1}^{k}[a_{i},b_{i}]$, which is insufficient to invoke [5, Lemma 2].
@@
By the same proof as in [10, Theorem 2.1], it follows that $Q$ is a quasi-copula. It remains to prove that it is $k$-increasing. By Lemma 3.4, it suffices to prove that $Q$ is $k$-increasing on each $k$-face of each box ${\mathcal{B}}_{\mathbb{I}}$ for $\mathbb{I}\in\{0,1\}^{d}$. By construction, this is true for $\mathbb{I}_{\text{main}}=(1,1,\ldots,1)$. Let us now consider $\mathbb{I}\in\{0,1\}^{d}\setminus\mathbb{I}_{\text{main}}$. If the face has constant 0 on some component, then $Q$ is constantly equal to 0 on the whole face and $Q$ is $k$-increasing. Otherwise every constant component is equal to $a_{j}$ or $1$. Further on, we can assume that all non-constant components are $a_{j},1]$ and not $0,a_{j}]$, because otherwise we can further project the section on the $j$-th component to $\{a_{j}\}$, hence obtaining a lower dimensional face of ${\mathcal{B}}_{\mathbb{I}_{\text{main}}}$. However, this reduction give us a $k$-dimensional face of ${\mathcal{B}}_{\mathbb{I}_{\text{main}}}$, whence it is $k$-increasing by construction. ∎
@@
The proof is the same as for [10, Proposition 3], but we include it for the sake of completeness. Let $S_{d}$ be the set of all permutations of a $d$-element set $\{1,\ldots,d\}$. For $\Phi\in S_{d}$ and $\mathbb{I}^{(j)}:=(\mathbb{I}_{1}^{(j)},\ldots,\mathbb{I}_{d}^{(j)})\in\{0,1\}^{d}$, let $\Phi(\mathbb{I}^{(j)}):=\big(\mathbb{I}_{\Phi(1)}^{(j)},\ldots,\mathbb{I}_{\Phi(d)}^{(j)}\big)$. For every optimal solution $\big(a_{1}^{\ast},\ldots,a_{d}^{\ast},b_{1}^{\ast},\ldots,b_{d}^{\ast},q_{\mathbb{I}^{(1)}}^{\ast},\ldots,q_{\mathbb{I}^{(2^{d})}}^{\ast}\big)$ to (3.2), $\big(a_{1}^{\ast},\ldots,a_{d}^{\ast},b_{1}^{\ast},\ldots,b_{d}^{\ast},q_{\mathbb{I}^{(1)}}^{\ast},\ldots,q_{\mathbb{I}^{(2^{d})}}^{\ast}\big)$ is also an optimal solution to (3.2), whence an optimal solution to (3.2) of the form (3.7) is equal to $\frac{1}{d!}\sum_{\Phi\in S_{d}}\big(a_{\Phi(1)}^{\ast},\ldots,a_{\Phi(d)}^{\ast},b_{\Phi(1)}^{\ast},\ldots,b_{\Phi(d)}^{\ast},q_{\Phi(\mathbb{I}^{(1)})}^{\ast},\ldots,q_{\Phi(\mathbb{I}^{(2^{d})})}^{\ast}\big).$
@@
We start by $k=2$. By [11, Proposition 5], (3.8) is equivalent to